OpenStudy (anonymous):

Carla runs 4 miles every day. She jogs from home to the school track, which is 3/4 mile away. She then runs laps around the 1/4 mile track. Carla then jogs home. How many laps does she run at school?

OpenStudy (anonymous):

4 - 3/4-3/4=4-1.5=2.5 /(1/4)=10laps

OpenStudy (anonymous):

thanks so much you are a lifesaver

OpenStudy (anonymous):

good answer